# Semantic Faithfulness and Entropy Production Measures to Tame Your LLM Demons and Manage Hallucinations

### v1

@@ 5.5 Qualitative Evaluation with LLM-as-a-Judge
Our hypothesis is that the Semantic Faithfulness metric $\mathcal{F}_{S}$ and Semantic Entropy Production (SEP) provide effective guidance for answer selection, capturing dimensions of quality that human evaluators would recognize: structural alignment with question requirements, comprehensive coverage of requested topics, appropriate grounding in context, and absence of tangential or repetitive content.
We compare answers with the highest and lowest $\mathcal{F}_{S}$ scores from a separate experimental run using Group A comprehensive risk questions. [^4] The $\mathcal{F}_{S}$ values reported in this section (0.872 and 0.250) differ from those in Table 1 because they were computed in a different experimental session with different random initialization. This variability is expected and does not affect the qualitative conclusions. For each pair, we present both answers to the LLM with instructions to analyze faithfulness, completeness, coherence, and relevance. The qualitative analysis reported below is provided verbatim as generated by Claude Sonnet 4.5.
@@
Answer A ($\mathcal{F}_{S}=0.872$) produced a structured response organized around four numbered sections: (1) Supply Chain Dependencies and Demand Forecasting Challenges, (2) Competitive Risks and Operational Challenges, (3) Impact of External Factors: Export Controls and Macroeconomics, and (4) Financial Illustration and Customer Concentration. The answer maintains direct alignment with the question’s framework and concludes with a concise summary paragraph that ties all elements together.
Answer B ($\mathcal{F}_{S}=0.250$) produced a longer response with an executive summary, more detailed subsections, and extensive bullet points. While covering similar material, the answer includes additional elements such as cybersecurity threats and EU AI Act regulatory compliance, and provides a more extensive final section on “Implications for Financial Health and Long-Term Standing.”
@@
The judge’s detailed explanation noted: “Both answers provide excellent, comprehensive assessments of NVIDIA’s vulnerabilities that are highly faithful to the context.” Key observations include:
  • Both answers accurately describe supply chain dependencies on third-party manufacturers in geopolitically sensitive regions (Taiwan, China)
  • Both explain the long lead times (12+ months) and demand forecasting challenges
  • Both cite the same specific financial examples: inventory provisions for low-yielding Blackwell material (Q2 FY2025) and warranty liability from third-party component defects (FY2023)
  • Both discuss competitive risks from rivals and customers developing their own solutions
  • Both address customer concentration and export controls comprehensively
The judge characterized the differences as “stylistic rather than substantive”: Answer A is “slightly more focused and concise,” while Answer B is “slightly more comprehensive in scope” with additional analytical framing.
@@
The LLM judge’s inability to distinguish between answers with substantially different $\mathcal{F}_{S}$ scores (0.872 vs 0.250) illustrates that these two evaluation approaches capture different aspects of answer quality. While this paper reports only one detailed example, additional experiments (not included here) show that LLM judges sometimes agree and sometimes disagree with $\mathcal{F}_{S}$-based rankings. Notably, we observed cases where $\mathcal{F}_{S}$ correctly identified inferior answers that the LLM judge rated highly, while also failing to detect hallucinated content.
Such absence of coherence between the two approaches is neither surprising nor problematic: $\mathcal{F}_{S}$ measures information-theoretic alignment between the answer’s semantic distribution and the optimal channel implied by the question-context pair, while LLM judges evaluate surface-level criteria such as coherence, completeness, and relevance. An answer can score well on traditional quality metrics while exhibiting suboptimal information-theoretic alignment—for instance, by including contextually accurate but tangential topics that dilute focus on the question’s core requirements.
We therefore recommend using $\mathcal{F}_{S}$ and LLM-based evaluation as complementary tools rather than expecting agreement. The Semantic Faithfulness metric provides a principled, quantitative measure of alignment with implicit information requirements that subjective evaluation may miss, particularly for subtle semantic drift or hallucination. Complete examples including question-answer pairs and precomputed distributions are available in the project repository’s data/cache/ directory.
@@ 6 Conclusions
LLM-as-a-Judge Validation.
Qualitative evaluation using Claude Sonnet 4.5 showed that LLM judges and $\mathcal{F}_{S}$ capture different aspects of answer quality. While judges evaluate surface-level criteria (coherence, completeness, relevance), $\mathcal{F}_{S}$ measures information-theoretic alignment. Additional experiments beyond those reported here revealed cases where the two approaches disagreed—including instances where $\mathcal{F}_{S}$ correctly identified inferior answers that the LLM judge rated highly due to missed hallucinations. We recommend using both as complementary evaluation tools.

### v2

@@ 5.5 Qualitative Evaluation with LLM-as-a-Judge
Our hypothesis is that the Semantic Faithfulness metric $\mathcal{F}_{S}$ and Semantic Entropy Production (SEP) provide effective guidance for answer selection, capturing dimensions of quality that human evaluators would recognize: structural alignment with question requirements, comprehensive coverage of requested topics, appropriate grounding in context, and absence of tangential or hallucinated content.
We compare answers with the highest and lowest $\mathcal{F}_{S}$ scores from a separate experimental run using Group A comprehensive risk questions. [^4] The $\mathcal{F}_{S}$ values reported in this section (0.324 and 0.250) differ from those in Table 1 because they were computed in a different experimental session with different random initialization. This variability is expected and does not affect the qualitative conclusions. For each pair, we present both answers to the LLM judge with instructions to analyze faithfulness, completeness, coherence, and relevance. The qualitative analysis reported below is provided verbatim as generated by Claude Sonnet 4.5.
@@
Answer A ($\mathcal{F}_{S}=0.324$) produced a structured response organized around the question’s framework with sections on: Vulnerabilities in Manufacturing and Supply Networks, Competitive Threats to Market Leadership, Organizational Capacity for Navigating Technology Transitions, and Impact of Broader Forces on Operational Risks. The answer maintains direct alignment with the question’s structure, uses precise language from the source context, and correctly states that “three direct customers representing 12%, 11%, and 11% of total revenue in fiscal year 2025.”
Answer B ($\mathcal{F}_{S}=0.250$) produced a longer response with an executive summary, numbered sections, and extensive bullet points. While covering similar material, the answer includes additional elements such as cybersecurity threats and provides a more extensive final section on “Implications for Financial Health and Long-Term Standing.” Critically, Answer B contains a subtle hallucination: it states that “Customers A, B, and C represented 12%, 11%, and 11% of total revenue”. By this, it fabricates specific customer names that do not appear in the source context, which only mentions “three direct customers” without naming them.
@@
The judge’s detailed explanation noted: “Both answers provide comprehensive, well-structured analyses of NVIDIA’s risk landscape that are highly faithful to the context.” Remarkably, the LLM judge identified the phrase “Customers A, B, and C represented 12%, 11%, and 11% of total revenue” as a “distinctive phrase” in Answer B, yet failed to recognize it as fabricated content. The judge concluded that “Neither contains significant hallucinations or omissions. The differences are primarily stylistic rather than substantive, making this effectively a tie in quality.”
The SF Score Disfavors Hallucinated Content.
This example illustrates that the LLM judge may overlook (and even praise!) hallucinated content. The source document states only that NVIDIA has “three direct customers” representing the cited revenue percentages, but it never names these customers. Answer B invented “Customers A, B, and C” as placeholder names, a subtle but consequential hallucination that could mislead downstream systems or users into researching non-existent entities.
On the other hand, the $\mathcal{F}_{S}$ metric correctly assigned a lower faithfulness score (0.250) to the hallucinating answer compared to the faithful answer (0.324), capturing semantic deviation that the LLM judge’s surface-level evaluation missed entirely.
@@
The LLM judge’s inability to detect fabricated content that $\mathcal{F}_{S}$ correctly penalized demonstrates that these two evaluation approaches capture fundamentally different aspects of answer quality. While this paper reports one detailed example, additional experiments (not included here) show that LLM judges sometimes agree and sometimes disagree with $\mathcal{F}_{S}$-based rankings. [^5] Complete examples including question-answer pairs and precomputed distributions are available in the project repository’s data/cache/ and docs/examples/ directories.
This divergence is neither surprising nor problematic: $\mathcal{F}_{S}$ measures information-theoretic alignment between the answer’s semantic distribution and the optimal channel implied by the question-context pair, while LLM judges evaluate surface-level criteria such as coherence, completeness, and relevance. An answer can score well on traditional quality metrics while containing hallucinated content that sounds plausible but deviates from the source material.
We therefore recommend using $\mathcal{F}_{S}$ and LLM-based evaluation as complementary tools rather than expecting agreement. The Semantic Faithfulness metric provides a principled, quantitative measure of alignment with implicit information requirements that subjective evaluation may miss, particularly for subtle semantic drift or hallucination such as fabricated entity names.
@@ 6 Conclusions
Qualitative evaluation demonstrated that LLM judges and $\mathcal{F}_{S}$ capture fundamentally different aspects of answer quality. In our detailed example, the LLM judge rated two answers identically (9/10 across all criteria) despite one containing fabricated customer names (a hallucination the judge explicitly identified as a “distinctive phrase”), yet failed to recognize as fabricated content. The $\mathcal{F}_{S}$ metric correctly assigned a lower score to the hallucinating answer (0.250 vs 0.324), capturing semantic deviation that surface-level evaluation missed. This divergence reflects the complementary nature of the two approaches: LLM judges evaluate coherence, completeness, and relevance, while $\mathcal{F}_{S}$ measures information-theoretic alignment with the source context. We recommend using both as complementary evaluation tools, with $\mathcal{F}_{S}$ providing particular value for detecting subtle hallucinations that sound plausible but deviate from source material.